mw104969:

CAN SOMEONE PLLZZZ HELP ME!!!! 3x^3-20x^2+12x=0

umm:

 Does your question look like this?

mw104969:

yes

umm:

Because if so, here's step by step.. First things first, let's left side of equation. \[x(3x-2) (x-6)=0\] Second, we'll set factors equal to 0. \[x=0\]or \[3x-2=0\]or \[x-6=0\] Then we get your final answer. \[x=0\]or \[x=\frac{ 2 }{ 3 }\]or \[x=6\]

mw104969:

ty

umm:

Anytime.